Katrina's age is twice of Nico's .The sum of their ages is 36
Answer by ValorousDawn(53)

Let Katrina's age be K, and Nico's age be N.
Katrina's age is twice of Nico's. K=2N
The sum of their ages is 36. K+N=36.
You can use substitution to get 2N+N=36 3N=36 N=12
We plug it back in to get Katrina's age. K=2N K=2(12) K=24.

Katrina's 24, Nico's 12.